40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 42     Entry time: Wed Oct 31 23:55:17 2007
Author: waldman 
Type: Other 
Category: OMC 
Subject: QPD tests 
The 4 QPDs for the OMC have been installed in the 056 at the test setup. All 4 QPDs work and have medm screens located under C2TPT. The breadboard mounted QPDs are not very well centered so their signal is somewhat crappy. But all 4 QPDs definitely see plenty of light. I include light and dark spectra below. QPDs 1-2 are table-mounted and QPD 2 is labeled with a bit of blue tape. QDPs 3-4 are mounted on the OMC. QPD3 is the near field detector and QPD4 is the far field. In other words, QPD3 is closest to the input coupler and QPD4 is farthest.

Included below are some spectra of the QPDs with and without light. For QPDs 1 & 2, the light source is just room lights, while 3&4 have the laser in the nominal OMC configuration with a few mWs as source. The noise at 100 Hz is about 100 microvolts / rtHz. If I recall correctly, the QPDs have 5 kOhm transimpedance (right Rich?) so this is 20 nanoamps / rtHz of current noise at the QPD.
Attachment 1: QPD_SignalSpectrum.pdf  141 kB  | Hide | Hide all
QPD_SignalSpectrum.pdf
Attachment 2: QPD_SignalSpectrum.gif  60 kB  | Hide | Hide all
QPD_SignalSpectrum.gif
ELOG V3.1.3-